# Enhancing Material Screening at Boulby Underground Laboratory with XIA UltraLo-1800 Alpha Particle Detectors

### v1

@@ 1 Introduction
Among the 36 radon isotopes, ranging in atomic masses from 193 to 228, all are radioactive with the most stable,  [^2] Rn, constituting part of the 238U decay chain and exhibiting a half-life of 3.82 days [4]. The next most common isotopes are 220Rn and 219Rn, which belong to the 232Th and 235U decay chains, respectively [4]. Radon, due to its noble gas characteristics, resists chemical interactions, posing a challenge for its removal [7]. Furthermore, as a monatomic gas, it possesses a long diffusion length in solids [6, 8], which, combined with its half-life, results in a resistance to effective sealing. Decay of radon generates electrically charged daughter ions which, due to electrostatic attraction, ’plate out’ and adhere to nearby surfaces [9]. Subsequent decays of surface deposits may implant themselves to depths of several micrometres, thereby hindering removal efforts [9]. Several radon chain daughter isotopes, notably 210Po (138 days) and 210Pb (22 years), have long half-lives against alpha decay. Should these isotopes be in the vicinity of instrument materials with large ($\alpha$, n) cross sections, such as 19F found in PTFE, then their subsequent alpha decay can produce neutrons within the instrument [10]. These neutrons introduce a problematic background for experiments such as those searching for weakly interacting massive particles, where the signal being searched for is virtually identical to that produced by neutron single scatters.
@@
    Characterizing the XIA UltraLo-1800 detector’s performance, focusing on energy reconstruction and background radioactivity levels.
@@
    Analyzing the detector’s stability.
@@ 2.1 Detector Calibration and Energy Reconstruction
The Boulby UltraLo-1800 was routinely calibrated using a custom-made 210Po alpha-particle source provided by Spectrum Techniques [20]. 210Po decays by emitting a 5.408 MeV alpha particle with a branching ratio of essentially 100% [21]. To better understand the detector’s response to various materials and surface contamination, we constructed comprehensive simulations using the GEANT4-v10.5.1 toolkit [22]; G4EmStandardPhysics, PhysListEmPenelope, and PhysListEmLivermore were used to model electromagnetic processes, while the G4HadronElasticPhysicsHP was used to model hadronic interactions. The geometry of the XIA UltraLo-1800 detector was implemented in the simulation, including all physical aspects, such as the active volume, anode, and guard rail, as well as its immediate surroundings. The 210Po source was modelled as an isotropic source positioned at the centre of the sample tray.
@@ 2.2 Background Measurements
The detector used for this study has been fully operational since January 2018. The sample tray is a critical component of the detector, as it holds the materials being assayed and determines the sensitive area exposed to the detector. To understand the inherent background measurements, we started by conducting measurements with the manufacturer-supplied stainless steel sample tray empty. A photograph of the stainless steel sample tray is shown in the left-hand panel of Figure 3. Several measurements were made, shown as the filled red data points in Figure 4, indicating an average activity of 1.3 $\pm$ 0.04 $\alpha$/cm2/khr. This result is in good agreement with results from other UltraLo-1800 facilities [16, 17, 18].
@@
Our long-term study of the UltraLo-1800 with this PTFE liner inserted (but otherwise empty) is presented as the filled blue squares in Figure 4. The mean emissivity is measured to be 0.15 $\pm$ 0.01 $\alpha$/cm2/khr; individual measurements have a standard deviation of 0.01, indicating the stability of these measurements.
@@
Similar efforts to quantify the absolute detector background have been reported by IBM [24] and XMASS [25], resulting in 0.6 $\pm$ 0.1 $\alpha$/cm2/khr and 0.14 $\pm$ 0.03 $\alpha$/cm2/khr, respectively. The detector’s operation near sea level may influence the IBM result, where cosmic radiation events contribute to the observed event rate. The XMASS detector, located underground at the Kamioka laboratory, produced results that are in statistical agreement with those reported here. The residual event rate observed in both the Boulby and Kamioka data is likely due to a combination of alpha emissions from detector construction materials and radon present in the argon buffer gas.
@@ 3 Surface and Bulk Alpha Measurements
Materials exposed to environments containing radon develop plate-out on their surfaces [9]. The dominant source of alpha-particle emission from high-purity material surfaces in environments with radon exposure is 210Po, due to its accumulation from radon progeny over the material’s lifetime. The other decay progeny, a radionuclide with a significant half-life, 210Pb, decays almost entirely by beta particle emission [4]. Consequently, unless exceptional care has been taken, most detector components manufactured in the last few decades will contain 210Po mixed into the bulk during production and manufacturing. By "bulk," we mean the internal substance of the material, which in some cases may include high surface area bulk products like powders for PTFE.
Alpha particle counting for samples often shows two primary components: 5.408 MeV alpha particles emitted by surface-deposited 210Po and lower energy alpha particles emitted from 210Po embedded within the bulk. The latter has lost some of its energy traversing to the material surface. Depending on the material, alpha particles from the bulk can be detected from various depths, up to several tens of micrometres in solids.
@@
The left-hand panel predicts the alpha particle yield detected from emission throughout layers of sequentially increasing depth, illustrating how emission from different depths contributes to the overall rate. Regarding bulk emission, the efficiencies noted in the left-hand figure illustrate how decay depth impacts the probability of an alpha particle escaping the bulk. Decays occurring near the surface have almost an equal chance of being emitted or buried within the sample. However, decays occurring at greater depths have progressively smaller chances of escaping the material.
@@
During sample measurements, additional alpha particle signals are expected due to the decay of radioisotopes present in the detector’s argon gas. This would generate broadly Gaussian features at 6.115 MeV (from decays of 218Po), 6.207 MeV (212Bi), 6.906 MeV (216Po), 7.834 MeV (214Po) and 8.954 MeV (212Po) [21]. The detected yields for each of these contributions reflect the activity level of the radionuclide in the gas, its decay branching ratio, and the detector efficiency at that energy [19]. Absolute detector efficiencies were based on data from XIA, validated at 5.408 MeV using 210Po calibration data, and extrapolated as required to other energies.
@@ 3.1 Assay of Copper Plate-out and Bulk Activity
The copper sheet was then placed in the BUGS cleanroom, where it was exposed to 2.30 $\pm$ 0.03 Bq/m3 of radon for 458 days [26]. This exposure duration was chosen based on our prior experience with radon exposure and aimed to achieve significant plate-out of radon progeny, including 210Po, onto the copper surface, thus increasing the activity levels.
@@
Table 1: The measured activities of 210Po emanating from an ultra-pure copper sheet along with implications of each state for cleanliness procedures.
@@ 4 Conclusion
The XIA UltraLo-1800 detector, installed within the Boulby UnderGround Screening (BUGS) facility, has been characterized and shown to maintain stable internal backgrounds over extended periods. This stability is essential in discerning minute levels of radioactivity in materials used for rare-event searches. A GEANT4 simulation model was developed to effectively differentiate between bulk and surface alpha emissions, a distinction vital for understanding and mitigating contamination sources in detector materials. Specifically, our model simulates alpha particles’ behaviour emitted from both the surface and the bulk, providing insights into how their emission depth affects detection probabilities.
Measurements carried out on ultra-pure electroformed copper have demonstrated a substantial improvement of up to threefold compared to previous measurements obtained from PTFE. This significant enhancement directly results from operating within a deep underground laboratory, which effectively shields the detector from cosmic ray-induced backgrounds. Furthermore, the UltraLo-1800 detector has been employed to quantify the effectiveness of different surface-cleaning techniques for materials commonly used in particle physics.
In this work, we have demonstrated the UltraLo-1800 detector’s capability to precisely measure both surface and bulk alpha emissions, confirming the predictions of our GEANT4 simulation model. This capability makes the UltraLo-1800 an essential component of the BUGS facility’s material screening program.

### v3

@@ 1 Introduction
Among the 36 radon isotopes, ranging in atomic masses from 193 to 228, all are radioactive with the most stable,  [^2] Rn, constituting part of the 238U decay chain and exhibiting a half-life of 3.82 days [4]. The next most common isotopes are 220Rn and 219Rn, which belong to the 232Th and 235U decay chains, respectively [4]. Radon, due to its noble gas characteristics, resists chemical interactions, posing a challenge for its removal [7]. Furthermore, as a monatomic gas, it possesses a long diffusion length in solids [6, 8], which, combined with its half-life, results in a resistance to effective sealing. Decay of radon generates electrically charged daughter ions which, due to electrostatic attraction, ‘plate out’ and adhere to nearby surfaces [9]. Subsequent decays of surface deposits may implant themselves to depths of several micrometres, thereby hindering removal efforts [9]. Several radon chain daughter isotopes, notably 210Po (138 days) and 210Pb (22 years), have long half-lives against alpha decay. Should these isotopes be in the vicinity of instrument materials with large ($\alpha$, n) cross sections, such as 19F found in PTFE, then their subsequent alpha decay can produce neutrons within the instrument [10]. These neutrons introduce a problematic background for experiments such as those searching for weakly interacting massive particles, where the signal being searched for is virtually identical to that produced by neutron single scatters.
@@
    Characterising the XIA UltraLo-1800 detector’s performance, focusing on energy reconstruction and background radioactivity levels.
@@
    Analysing the detector’s stability.
@@ 2.1 Detector Calibration and Energy Reconstruction
The Boulby UltraLo-1800 was routinely calibrated using a custom-made 210Po alpha-particle source provided by Spectrum Techniques [20]. 210Po decays by emitting a 5.304 MeV alpha particle with a branching ratio of essentially 100% [21]. To better understand the detector’s response to various materials and surface contamination, we constructed comprehensive simulations using the GEANT4-v10.5.1 toolkit [22]; G4EmStandardPhysics, PhysListEmPenelope, and PhysListEmLivermore were used to model electromagnetic processes, while the G4HadronElasticPhysicsHP was used to model hadronic interactions [23, 24]. The geometry of the XIA UltraLo-1800 detector was implemented in the simulation, including all physical aspects, such as the active volume, anode, and guard rail, as well as its immediate surroundings. The 210Po source was modelled as an isotropic source positioned at the centre of the sample tray.
@@ 2.2 Background Measurements
The detector used for this study has been fully operational since January 2018. The sample tray is a critical component of the detector, as it holds the materials being assayed and determines the sensitive area exposed to the detector. To understand the inherent background measurements, we started by conducting measurements with the manufacturer-supplied stainless steel sample tray empty. A photograph of the stainless steel sample tray is shown in the left-hand panel of Figure 3. Several measurements were made, shown as the filled red data points in Figure 4, indicating an average activity of 1.28 $\pm$ 0.03 $\alpha$/cm2/khr. This result is in good agreement with results from other UltraLo-1800 facilities [16, 17, 18].
@@
Our long-term study of the UltraLo-1800 with this PTFE liner inserted (but otherwise empty) is presented as the filled blue squares in Figure 4. The mean emissivity is measured to be 0.14 $\pm$ 0.01 $\alpha$/cm2/khr; individual measurements have a standard deviation of 0.01, indicating the stability of these measurements.
@@
Similar efforts to quantify the absolute detector background have been reported by IBM [26] and XMASS [27], resulting in 0.6 $\pm$ 0.1 $\alpha$/cm2/khr and 0.14 $\pm$ 0.03 $\alpha$/cm2/khr, respectively. The detector’s operation near sea level may influence the IBM result, where cosmic radiation events contribute to the observed event rate. The XMASS detector, located underground at the Kamioka laboratory, produced results that are in statistical agreement with those reported here. The residual event rate observed in both the Boulby and Kamioka data is likely due to a combination of alpha emissions from detector construction materials and radon present in the argon buffer gas. Despite our best efforts, some background levels are irreducible due to the detector’s material properties and operational requirements. We are exploring additional measures, such as argon scrubbing, to reduce this background further.
@@ 3 Surface and Bulk Alpha Measurements
Materials exposed to environments containing radon develop plate-out on their surfaces [9]. The dominant source of alpha-particle emission from high-purity material surfaces in environments with radon exposure is 210Po, due to its accumulation from radon progeny over the material’s lifetime. The other decay progeny, a radionuclide with a significant half-life, 210Pb, decays almost entirely by beta particle emission [4]. Consequently, unless exceptional care has been taken, most detector components manufactured in the last few decades will contain 210Po mixed into the bulk during production and manufacturing. By "bulk", we mean the internal substance of the material, which in some cases may include high surface area bulk products like powders for PTFE.
Alpha particle counting for samples often shows two primary components: 5.304 MeV alpha particles emitted by surface-deposited 210Po and lower energy alpha particles emitted from 210Po embedded within the bulk [28]. The latter has lost some of its energy traversing to the material surface. Depending on the material, alpha particles from the bulk can be detected from various depths, up to several tens of micrometres in solids.
In our simulations using the GEANT4 framework, we modelled the copper sample with 210Po uniformly distributed throughout the bulk up to a depth of 12 micrometres from the surface. This depth corresponds to the maximum range of 5.3 MeV alpha particles in copper, beyond which alpha particles emitted in bulk would not have sufficient energy to escape the material and be detected by the UltraLo-1800 detector. The activity concentration of 210Po was assumed constant within this depth, reflecting uniform bulk contamination.
Each depth layer contributes to the overall spectrum based on its volume and the uniform activity concentration. The simulation accounts for energy losses of alpha particles as they traverse the copper matrix, leading to the continuous energy distribution observed for bulk emissions.
@@
The left-hand panel of Figure 6 predicts the alpha particle yield detected from emission throughout layers of sequentially increasing depth, illustrating how emission from different depths contributes to the overall rate. Regarding bulk emission, the efficiencies noted in the left-hand figure illustrate how decay depth impacts the probability of an alpha particle escaping the bulk. Decays occurring near the surface have almost an equal chance of being emitted or buried within the sample. However, decays occurring at greater depths have progressively smaller chances of escaping the material.
@@
During sample measurements, additional alpha particle signals are expected due to the decay of radioisotopes present in the detector’s argon gas. This would generate broadly Gaussian features at 6.115 MeV (from decays of 218Po), 6.207 MeV (212Bi), 6.906 MeV (216Po), 7.834 MeV (214Po) and 8.954 MeV (212Po) [21], note that these values represent the total decay energies (Q-values) for each decay, and the alpha particles emitted carry slightly less energy due to the recoil of the daughter nucleus. The detected yields for each of these contributions reflect the activity level of the radionuclide in the gas, its decay branching ratio, and the detector efficiency at that energy [19]. Absolute detector efficiencies were based on data from XIA, validated at 5.304 MeV using 210Po calibration data, and extrapolated as required to other energies.
@@ 3.1 Assay of Copper Plate-out and Bulk Activity
The copper sheet was then placed in the BUGS cleanroom, where it was exposed to 2.30 $\pm$ 0.03 Bq/m3 of radon for 458 days. This exposure duration was chosen based on our prior experience with radon exposure and aimed to achieve significant plate-out of radon progeny, including 210Po, onto the copper surface, thus increasing the activity levels.
@@
Table 1: The measured activities of 210Po emanating from a copper sheet along with implications of each state for cleanliness procedures.
@@ 4 Conclusion
The XIA UltraLo-1800 detector, installed within the Boulby UnderGround Screening (BUGS) facility, has been characterised and shown to maintain stable internal backgrounds over extended periods. This stability is essential for detecting minute levels of radioactivity in materials used for rare-event searches. A GEANT4 simulation model was developed to effectively differentiate between bulk and surface alpha emissions, a distinction critical for understanding and mitigating contamination sources in detector materials. Specifically, our model simulates the behaviour of alpha particles emitted from both the surface and bulk, providing insights into how emission depth affects detection probabilities.
Based on the study of background rates from a variety of trays used in the XIA, we concluded that the PTFE-lined tray reduced the background rate by a factor of 9 compared to the default XIA stainless steel (SS) tray. However, further measurements showed that the ultra-pure electroformed copper tray produced the lowest background rate, and therefore the plan is to switch to the electroformed copper tray for future measurements. The reduced background rate is also directly attributable to operating within a deep underground laboratory, which effectively shields the detector from cosmic ray-induced backgrounds. Moreover, the UltraLo-1800 detector has been employed to quantify the effectiveness of various surface-cleaning techniques for materials commonly used in particle physics, such as copper. Through this work, we have identified more effective methods for reducing surface contamination, particularly crucial for mitigating radon daughter plate-out.
In this work, we have demonstrated the UltraLo-1800 detector’s capability to measure both surface and bulk alpha emissions, aided by our GEANT4 simulation model. This capability establishes the UltraLo-1800 as an essential component of the BUGS facility’s material screening program. These results have significant implications for future rare-event search experiments, such as BUTTON (Boulby Underground Technology Testbed for Observing Neutrinos) [31] and XLZD [32]. The advancements presented in this study pave the way for achieving the required levels of material purity and background suppression in these next-generation experiments.
Ongoing efforts to enhance BUGS’s material screening capabilities include the introduction of inductively coupled plasma mass spectrometry (ICP-MS) and a radon emanation detection system. Additionally, plans are underway to integrate in-house production of electroformed copper, which will further reduce background contamination and improve the sensitivity of future assays. These enhancements will solidify BUGS as a leading facility in underground science and rare-event detection, supporting the stringent demands of upcoming experiments.